Explain three different approaches to calculate the GDP. Which approach should be used in the economy like Pakistan and Why? (2 Marks, Maximum 200 words)


An economy produces two goods: tomatoes and ketchup. It is assumed that half of the tomatoes are bought and consumed as final good; the other half is used to produce ketchup.


In 2014, 30 KGs of tomatoes are produced at Rs.60 each, and 20 KGs of ketchup are produced at Rs.200 each.


In 2015, 40 KGs of tomatoes are produced at Rs.80 each, and 30 KGs of ketchup are produced at Rs.195 each.


Find the following:


Assuming 2014 is the base year; calculate the real, nominal GDP and GDP deflator in 2014 and 2015. (3 Marks)


1
Expert's answer
2020-11-27T13:15:07-0500

GDP can be measured in three different ways:


Income approach


The income method starts with the income earned from the production of goods and services. Under this approach, we calculate the income earned from all factors of production in an economy. Factors of production are the inputs that go into the production of the final product or service eg land, labor, and capital.


Formula:GDP=wages+rent+interest+profits+depreciation+net foreign factor income


Expenditure approach


The expenditure approach measures total expenditure incurred by all entities no goods and services within the domestic borders of a country.


formula:GDP=consumption expenditure+investment expenditure+Government expenditure+(Exports-Imports).


Production approach

this method measures the market value of all goods and services produced within the borders of a country. Real GDP is computed to avoid a distorted measure of GDP due to changes in prices.


formula:GDP=real GDP(at constant prices)-taxes +subsidies.


The Pakistan economy is faced with several challenges of a very wide budget and trade deficit, galloping inflation, and an increase in the level of poverty. Due to this reason, the most suitable approach in calculating its GDP is the expenditure method.


Q2. Answer:


Real GDP=nominal GDP/GDP deflator


Nominal GDP in 2014=(30*60)+(20*200)=5800

Therefore; real GDP=5800/100=58

The GDP deflator is always 100% in the base year.


Nominal GDP in 2015=(40*80)+(30*195)=9050

GDP deflator=9050/5800*100=156%

Therefore; real GDP=9050/156=58